OpenStudy (anonymous):

Is any of these numbers prime and which ones: 913, 1913, 2913, 3913, 4913, 5913, 6913, 7913, 8913, 9913? List them please! Just the primes.

OpenStudy (cwrw238):

you can eliminate some which are divisible by 3 add up digits and if they come to 3,6 or 9 they are dviisible by 3 eg 2913 = 2+9+1+3 = 15; 1+5 = 6 so 2913 is not prime

OpenStudy (anonymous):

the quickest way to check if a number is prime is by attempting to divide the number by other prime numbers starting with 2 up until you arrive at the square root of the number you are testing (rounded up). so for 913 you try dividing it by every prime number between 2 and \(\sqrt{913}\) (which is 31, rounded up) here's the first few prime numbers 2,3,5,7,11,13,17,19,23,29,31,37,41,43,47,53,59,61,67,71

OpenStudy (anonymous):

and then there are indeed tricks like the one cwrw described.

ganeshie8 (ganeshie8):

913, 1913, 2913, 3913, 4913, 5913, 6913, 7913, 8913, 9913 after divisibility by 3 check : 913, 1913, 3913, 4913, 6913, 7913, 9913 after divisibility by 7 check : 913, 1913, 4913, 6913, 7913, 9913 after divisibility by 11 check : 1913, 4913, 6913, 7913, 9913

ganeshie8 (ganeshie8):

None of these remaining are easy to figure out